OpenStudy (anonymous):

who can help me with this 5=14 sin x

OpenStudy (sparrow2):

x=arcsin5/14+pk , k is whole number p=3,14

OpenStudy (anonymous):

No i mean can you finish it up

OpenStudy (anonymous):



OpenStudy (sparrow2):

x=arcsin5/15

OpenStudy (anonymous):

can u please break it down

OpenStudy (anonymous):

who can help me with this

myininaya (myininaya):

What question are you asking? Are you asking to solve 5=14sin(x)?

OpenStudy (anonymous):

i just need someone to finish it up

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

do i have to divide through by 14sin

OpenStudy (sparrow2):

oh no :)

myininaya (myininaya):

You need to first isolate the sin(x) by dividing 14 on both sides

OpenStudy (anonymous):

okay then

myininaya (myininaya):

taking arcsin( ) of both sides will give you a solution in the interval (-pi/2,pi/2)

OpenStudy (anonymous):

20.92

OpenStudy (anonymous):

is that correct

myininaya (myininaya):

If you are looking for a solution in degrees yes that is one solution to the equation posted there are infinitely many more though

OpenStudy (anonymous):

i approximated my answer

OpenStudy (anonymous):

am just finding x

myininaya (myininaya):

Are you looking for all the solutions?

myininaya (myininaya):

are you solving in a a certain interval?

myininaya (myininaya):

are you allowed to approximate?

OpenStudy (anonymous):

The question has no instruction about approximation it only says find x

OpenStudy (anonymous):

but am i correct

myininaya (myininaya):

\[\sin(x)=a \\ \text{ where } a \in (-1,1) \\ \\ \text{ can be solved by using the following } \\ \sin(x+2 n \pi)=a \\ \text{ also } -\sin(x+\pi+2n \pi)=a \\ \text{ take solving } \sin(x)=\frac{-1}{2} \\ \implies \sin(x+2n \pi)=\frac{-1}{2} \implies x+ 2 n \pi=\arcsin(\frac{-1}{2}) \\ \implies x=\arcsin(\frac{-1}{2})-2n \pi \\ -\sin(x+\pi+2n \pi)=- \frac{1}{2} \implies \sin(x+\pi+2 n \pi)=\frac{1}{2} \implies x+\pi+2 n \pi= \arcsin(\frac{1}{2}) \\ \implies x=\arcsin(\frac{1}{2})-\pi-2 \pi n\]